OpenStudy (ofmiceparade):

g(x) = 3x2 + 12x + 15 HOW DO I FIND THE AXIS OF SYMMETRY

OpenStudy (ofmiceparade):

@Vampirelove28

OpenStudy (jaden20):

This might help? http://www.tiger-algebra.com/drill/3X~2-12x-15=0/

OpenStudy (ofmiceparade):

I think maybe?

OpenStudy (ofmiceparade):

I have to explain why though

OpenStudy (anonymous):

To find the vertex, convert to vertex form by completing the square. 3x^2-12x+15 = 3 (x^2 - 4x) + 15 = 3 (x^2 - 4x + 4 - 4) + 15 = 3 {(x^2 - 2x - 2x + 4) - 4} + 15 = 3 {x(x-2) -2(x-2)} - 12 + 15 = 3 (x - 2)^2 + 3 Original vertex form is a (x - h)^2 + k where (h, k) are the co-ordinates of the vertex. Therefore, the vertex is (2, 3). Also, since the y co-ordinate is 3, the axis of symmetry is x = 3. and the minimum value of the quadratic equation is y = 3. [Because a > 0 the parabola is opening upwards and there is a minimum value] Hope this helps! :)

OpenStudy (jaden20):

Ya i think this guy has it right just use what he has...

OpenStudy (jaden20):

Ya i think this guy has it right just use what he has...